OpenStudy (anonymous):

wx^2+wl/2-wx simplified to wlx/2-wx^2/2 can someone show me how they got there thanks!

OpenStudy (anonymous):

is it : \[wx^2 - (wl/2)\]?

OpenStudy (anonymous):

wx^2 - wl/2 - wx *

OpenStudy (anonymous):

is this the question?

OpenStudy (anonymous):

+wl/2

OpenStudy (anonymous):

lol okay ^^"

OpenStudy (anonymous):

is it : wx^2 + (wl/2-wx) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

are you sure of your answer? because I got in the end :\[wx^2 + wl\]

OpenStudy (anonymous):

could you show me how you got wx^2 + wl

OpenStudy (anonymous):

steps, first same denominator:\[= (wx^2(2-wx) + wl) /2-wx\]\[= (2wx^2 -w^2x^3 + wl)/2-wx\] then multiply by 2+wx up and down:\[= (2wx^2-w^2x^3+wl)(2+wx)/(2-wx)(2+wx)\]\[= (4wx^2-w^3x^4 + 2wl + w^2xl)/(2-wx)(2+wx)\] then factor by grouping :\[= [wx^2(4-w^2x^2) + wl(2 + wx)]/(2-wx)(2+wx)\]\[= [wx^2(2-wx)(2+wx) + wl(2+wx)]/(2-wx)(2+wx)\]\[= [(2+wx)(2-wx)(wx^2+wl)]/(2-wx)(2+wx)\] simplify (cancel the commons) : \[= wx^2 + wl\] ^_^ correct me if I'm wrong please .

OpenStudy (anonymous):

thank you!!!

OpenStudy (anonymous):

a pleasure :)

OpenStudy (anonymous):

do you use a program to type the formulas?

OpenStudy (anonymous):

no, click on equation on your reply box and type the formula there ^_^

OpenStudy (anonymous):

"equation" *